OpenStudy (anonymous):

4x^3-16x^2+12=0 I'm very confused, can someone help me solve the equation? :(

OpenStudy (anonymous):

Your first step is to divide everything by 4.

OpenStudy (anonymous):

Your next step would be to notice x = 1 solves the equation.

OpenStudy (anonymous):

So is it -4x+3x?

OpenStudy (anonymous):

No, you divide by 4: x^3 - 4x^2 + 3 = 0 x = 1 solves this equation so factor out this solution: (x - 1)(x^2 - 3x - 3) Can you do the next step?

OpenStudy (anonymous):

Is it supposely put equal to zero?

OpenStudy (anonymous):

yes (x - 1)(x^2 - 3x - 3) = 0, i thought that much was clear.

OpenStudy (anonymous):

Well it would be (x-1)=0 +1 +1 so x=1 , right?

OpenStudy (anonymous):

Yes, I already told you that you must notice x = 1 is a solution. However, it is a cubic, so there are 3 solutions to x, so you have to find the remaining 2.

OpenStudy (anonymous):

Is it 3 ?

OpenStudy (mathmate):

Use the quadratic formula to find the two remaining roots. In case you do not remember, for y=ax^2+by+c, x=(-b +/- sqrt(b^2-4ac) )/(2a) Use it on the remainder of the expression, namely (x^2 - 3x - 3). It will not be a rational root (i.e. involves square-roots).